OpenStudy (anonymous):

What are the intercepts of y = -0.75x + 3.0?

OpenStudy (solomonzelman):

\(\Large\color{blue}{ y = -0.75x + 3.0}\) compare to \(\Large\color{blue}{ y = mx + b}\) where ` b` is the `y-intercept`.

OpenStudy (solomonzelman):

To find the x-intercept, you would need to plug in `0` instead of `y` and solve for x.

OpenStudy (anonymous):

What does mx mean?

OpenStudy (solomonzelman):

m is just a number.

OpenStudy (solomonzelman):

mx means m × x

OpenStudy (anonymous):

ok thanks for your help,but out of (4, 0) and (0, 3) (2, 0) and (0, 6) (12, 0) and (0, 3) (6, 0) and (0, 1)....what would be right?

OpenStudy (solomonzelman):

looks like x-intercept is the first option point, and the second point is the y-intercept. What do you think is the ` b ` in your equation (looking at and comparing with ` y=m x + b` )?

OpenStudy (anonymous):

i keep getting (4, 0) and (0, 3)

OpenStudy (solomonzelman):

yes, correct !

OpenStudy (anonymous):

Thanks for your help!

OpenStudy (solomonzelman):

Anytime !

OpenStudy (anonymous):

How about y = 4x − 2?

OpenStudy (solomonzelman):

y = 4x -2 Y-intercept y = 4x -2 y = 4(0) -2 y=-2 X-intercept y = 4x -2 (0) = 4x -2 0=4x-2 0 \(\normalsize\color{blue}{ +2 }\) =4x-2 \(\normalsize\color{blue}{ +2 }\) 2=4x 2 \(\normalsize\color{blue}{ \div 4 }\) =4x \(\normalsize\color{blue}{ \div 4 }\) ½ =x